# Are LLMs Good Literature Review Writers? Evaluating the Literature Review Writing Ability of Large Language Models

### v1

@@ 1 Introduction
  • Second, we collect 1,106 literature reviews from 51 journals across 5 disciplines as the ground truth. We then design three tasks: generating references, writing an abstract, and writing a literature review based on a given topic.
  • Then, we evaluate the writing results of the LLMs from multiple aspects, including the hallucination rate in generated references, factual consistency with human-written results, and the comparison of semantic coverage.
  • Finally, we assess four LLMs using the proposed framework. By analyzing the experimental results, we find that hallucinated references remain an unavoidable issue for current LLMs. Additionally, the performance of LLMs in writing literature reviews varies across different academic disciplines.
@@ 2 Related Work
Recent studies have increasingly focused on using LLMs for writing literature reviews. For example,  Wang et al. (2024) proposed a literature survey writing process for LLMs, called AutoSurvey. By RAG retrieving the latest relevant paper incorporates real-time knowledge. In evaluation, Muti-LLM-as-judge is used to evaluate citation quality and content quality.
@@ 3.1 Dataset Construction
Then, the dataset $D$ is the article set from $K$ journals, $D=\{j_{0},...,j_{k},...,j_{K}\}$, the number of journals is 51, so $K=50$. For each journal $j_{k}$, there is a set of articles $j_{k}=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles, and each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{l},...,r_{L}\}$.
@@ 3.2 Task Design
  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the 10 most relevant studies to the research topic. Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page. In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
@@
  • Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a literature review according to the research topic provided in the title, keywords, and abstract. The literature survey should be about 1000 words long. LLMs also need to back up claims by citing previous studies (with a total of 10 citations in the literature survey). In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
@@ 3.3 Evaluation Metrics
For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2. Then, for each journal $j_{k}$, we can obtain the accuracy of generated references $Acc_{ref.}(j_{k})$ based on Eq 3. Subsequently, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(j_{k})$ across journals, as shown in Eq 4. Finally, we defined the reference accuracy metric $Acc$ as Eq 4. A high accuracy metric indicates a low hallucination rate.
@@
$Entail_{j_{k}}=\frac{1}{M}\sum_{i=0}^{M}{Entail_{i}}$ (8)
$Entail=\frac{1}{K}\sum_{k=0}^{K}{Entail_{j_{k}}}$ (9)
Additionally, we use commonly employed semantic similarity metrics to calculate the semantic coverage between the content generated by LLMs and human-written content. Specifically, we apply cosine similarity and the ROUGE metric for evaluation.
@@ 4.1 Experimental Settings
Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1106 articles, the dataset distribution across journals as shown in Figure 7.
LLMs Selection. We select four LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, and Llama-3.2-3B-Instruct.
For the generated reference evaluation, we use Semantic Scholar as the external database. Recently, LLMs-as-judges has become more common Chen et al. (2024); Zheng et al. (2023); Shangyu et al. (2024). So, we employ TRUE Honovich et al. (2022), along with GPT-4o as NLI models for evaluating factual consistency in context. To compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings.
@@ 4.2 Main Results
Results for Task 1. By analyzing the results in Table 1, we observe that Claude-3.5 achieves the highest accuracy score and title search rate, whereas Llama-3.2 records the lowest performance on both metrics for Task 1. Additionally, when evaluating the accuracy of LLM-generated references, treating the author item as correct if the first author is identified correctly results in a 3–5% increase in accuracy scores across all models. This finding highlights that accurately generating complete author lists remains a significant challenge for LLMs.
@@
Qwen-2.5-72B 19.81 12.36 17.64
Llama-3.2-3B 14.36 3.39 6.86
GPT-4o 30.51 21.36 24.77
Claude-3.5-Sonnet 63.99 51.44 55.40
@@
Qwen-2.5-72B 80.22 68.69 95.02 40.61 8.78 20.12
Llama-3.2-3B 79.28 61.90 92.14 40.35 8.96 20.52
GPT-4o 80.96 77.74 96.50 40.70 8.56 19.86
Claude-3.5-Sonnet 81.17 78.10 96.77 41.13 8.99 20.00
@@
$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ Acc. Acc.(first author)
Qwen-2.5-72B 38.39 28.89 33.64
Llama-3.2-3B 21.00 4.85 7.27
GPT-4o 60.24 50.69 54.24
Claude-3.5-Sonnet 66.51 58.89 62.90
Table 3: The experimental results of the four LLMs in Task 3. “$S_{t}$” refers to the title retrieval rate as defined in Eq 5.“Acc.” represents the accuracy defined in Eq 5. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
@@
Qwen-2.5-72B 7.18 12.57 13.24 12.13 17.31 17.88
Llama-3.2-3B 5.97 8.48 8.67 8.02 8.17 8.33
GPT-4o 8.66 18.78 19.50 18.33 28.81 29.63
Claude-3.5-Sonnet 16.05 24.34 25.21 21.15 33.25 34.41
@@
Figure 3: Task 1: Accuracy scores $\uparrow$ across different disciplines.
Figure 4: Task 2: NLI scores (TRUE)$\uparrow$ across different disciplines.
Figure 5: Task 3: Accuracy scores $\uparrow$ across different disciplines.
@@ 4.3 Analyze LLM-Generated References from Different Dimensions
In both Task 1 and Task 3, we ask LLMs to generate references. The overall performance was discussed in the previous section. In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2. As seen in Figure LABEL:f2a, Claude-3.5 demonstrates a clear advantage over other models across all dimensions in Task 1. Additionally, the accuracy of reference generation in Task 1 for Claude-3.5, GPT-4o, and Qwen-2.5 shows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension. The accuracy for journal name, page, and author is also relatively high. In contrast, Llama-3.2 shows higher accuracy in the page and author dimensions compared to other dimensions, but overall, Llama-3.2 does not exhibit a competitive advantage in reference generation accuracy.
Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5, GPT-4o, and Qwen-2.5. Notably, the accuracy of GPT-4o in certain dimensions approaches that of Claude-3.5. However, the performance of LLaMA-3.2 remains suboptimal.
@@ 4.4 Cross-Disciplinary Analysis
We then present bar charts in Figures  345, which illustrate the performance of different models across various tasks and disciplines.
First, we observe that in Task 1, as shown in Figure 3, almost all models exhibit the highest accuracy in the Mathematics discipline. Additionally, Claude-3.5 demonstrates relatively high accuracy in both Biology and Social science. GPT-4o, besides having the highest accuracy in Mathematics, shows higher accuracy in Physics and Social science compared to Biology and Technology. Qwen-2.5 also exhibits a relatively high accuracy in Physics, but it has the lowest accuracy in Chemistry. Llama-3.2 has the highest accuracy in Mathematics, but its accuracy in Technology is the lowest. Finally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences. The results are as follows: Claude-3.5(f=7.65, p<.001), GPT-4o(f=11.87, p<.001), Qwen-2.5(f=12.39, p<.001), Llama-3.2(f=10.78, p<.001). These results indicate significant differences in performance across disciplines for each model.
Secondly, as shown in Figure 4, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2. We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology. In terms of this metric, GPT-4o outperforms Claude-3.5 in Mathematics and Technology, indicating that while GPT-4o has a lower accuracy in reference generation, the difference between the two models in the abstract generation task is relatively small. Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines. The results are as follows: Llama-3.2 (f= 11.06, p<.001) Qwen2.5 (f= 16.88, p<.001) GPT-4o(f= 18.88, p<.001) Claude-3.5(f= 23.72, p<.001). These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
Finally, we examine the accuracy of each model across five disciplines in Task 3, as illustrated in Figure 4. It is evident that the accuracy of Claude-3.5 and GPT-4o are significantly higher than those of Qwen-2.5 and LLaMA-3.2 across all disciplines. Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology. The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5(f=33.33, p<.001), GPT-4o(f=39.09, p<.001), Qwen-2.5(f=32.84, p<.001), Llama-3.2(f=39.02, p<.001), indicating significant differences in each models’ performance across the various disciplines.
@@ 4.5 Comparison of LLM-Cited and Human-cited References
In this section, we also compare the LLM-generated references with those from the human-written original articles. Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article. As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5 are also cited in the human-written article, indicating a relatively high overlap rate. Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3. Notably, GPT-4o models show a 10% increase in the overlap rate. Next, we provide a more detailed comparison of the references across various dimensions. As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower. For Task 3, Claude-3.5 and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1. The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension.
@@ 5 Conclusion
We select four LLMs for task evaluation and find that Claude-3.5-Sonnet outperforms GPT-4o, Qwen-2.5-72B, and Llama-3.2-3B across all three tasks, particularly excelling in the task of generating accurate references. This advantage is likely influenced by the training data of each model. Additionally, we observed that each model has different strengths across disciplines. Overall, for the reference generation task, nearly all models perform better in Mathematics, while their performance is weaker in Chemistry and Technology. However, when writing abstracts, all models exhibit the lowest factual consistency in Social Science, as indicated by the entailment scores, compared to human-written texts.

### v2

@@ 1 Introduction
  • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth. We then design three tasks: generating references, writing an abstract, and writing a literature review based on a given topic.
  • Then, we evaluate the generated results of the LLMs from multiple aspects, including the hallucination rate in generated references, factual consistency with human-written results, and the comparison of semantic coverage.
  • Finally, we assess four LLMs using the proposed framework. By analyzing the experimental results, we find that hallucinated references remain an unavoidable issue for current LLMs. Additionally, the performance of LLMs in writing literature reviews varies across different disciplines.
@@ 2 Related Work
Recent studies have increasingly focused on using LLMs for writing literature reviews. For example,  Wang et al. (2024) proposed a literature survey writing process for LLMs, called AutoSurvey. By RAG retrieving the latest relevant paper incorporates real-time knowledge. In evaluation, Muti-LLM-as-judge is used to evaluate citation quality and content quality.  Wang_Hu_Wang_Yan_Sheng_He_2024 compared the use of LLMs across different academic tasks and found that researchers are more likely to use LLMs for literature review (LR) tasks than for paper understanding (PU) tasks.
@@ 3.1 Dataset Construction
Then, the dataset $D$ is the article set from $K$ journals, $D=\{j_{0},...,j_{k},...,j_{K}\}$, the number of journals is 51, so $K=50$. For each journal $j_{k}$, there is a set of articles $j_{k}=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles in $j_{k}$, and each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{l},...,r_{L}\}$.
@@ 3.2 Task Design
  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies to the research topic. Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page. In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
@@
  • Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a literature review according to the research topic provided in the title, keywords, and abstract. The literature survey should be about 1000 words long. LLMs also need to back up claims by citing previous studies (with a total of $N$ citations in the literature survey). In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
@@ 3.3 Evaluation Metrics
For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2. Then, for each journal $j_{k}$, we can obtain the accuracy of generated references $Acc_{ref.}(j_{k})$ based on Eq 3. Subsequently, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(j_{k})$ across journals, as shown in Eq 4. Finally, we defined the reference accuracy metric $Acc$ as Eq 4. A higher accuracy metric indicates a lower hallucination rate.
@@
$Entail=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}{Entail_{i}}$ (8)
Additionally, we use commonly employed semantic similarity metrics and Key Point Recall (KPR) to calculate the semantic coverage between the context generated by LLMs and human-written context. Specifically, for Task 2, we apply cosine similarity and the ROUGE metric for semantic coverage evaluation. For Task 3, we use the ROUGE metric and KPR to measure the semantic coverage of the literature review generated by the LLMs relative to human-written content.
KPR, first proposed by  Qi_Xu_Guo_Wang_Zhang_Xu_2024, is a metric designed to evaluate the effectiveness of LLMs in utilizing RAG for long documents. Since the human-written literature review is lengthy and difficult to compare directly, we adapt the KPR metric by first using GPT-4 to extract $q$ key points $X_{i}=[x_{i_{1}},x_{i_{2}},...,x_{i_{q}}]$ from the human-written literature review $c_{i}$, and then calculating the coverage of these key points by the model-generated literature review as Eq. 9.
$KPR=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (9)
where $c_{i}^{g}$ denotes the literature review generated by LLMs. Finally, we also concatenate key points and compute the ROUGE metric between the key point and $c_{i}^{g}$.
@@ 4.1 Experimental Settings
Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1105 articles, the dataset distribution across journals as shown in Appendix B Figure 4.
LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3 and Llama-3.2-3B-Instruct.
In task 1 and task 3, we set $N$ as 10, each model generates 10 references. For the generated reference evaluation, we use Semantic Scholar as the external database. Recently, LLMs-as-judges has become more common Chen et al. (2024); Zheng et al. (2023); Shangyu et al. (2024). So, for Task 2, we employ TRUE Honovich et al. (2022), along with GPT-4o as NLI models for evaluating factual consistency in context; To compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings. For task 3, we employ GPT-4o as the NLI model in Eq 9, and set $q$ as 10.
@@ 4.2 Main Results
Results for Task 1. By analyzing the results in Table 1, we observe that Claude-3.5 achieves the highest accuracy and title search rate, whereas Llama-3.2 exhibits the lowest performance on both metrics for Task 1. Furthermore, when evaluating the accuracy of LLM-generated references, treating the author item as correct when the first author is identified correctly results in a 3–5% increase in accuracy scores across all models. This finding underscores that accurately generating complete author lists remains a significant challenge for LLMs. Additionally, we conduct a year-wise analysis of the literature that each model cited correctly, as illustrated in the line graph in Figure 2. The graph shows that the majority of the correctly cited references are concentrated between 2010 and 2020.
@@
On the other hand, we find that Claude-3.5 achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5 recalls the most claims from the human-written literature review. While the literature review generated by DeepSeek-V3 performs well on the ROUGE metrics.
@@
Qwen-2.5-72B 21.80 12.25 17.58
Llama-3.2-3B 16.62 3.45 6.95
DeepSeek-V3 56.04 46.33 50.66
GPT-4o 32.07 21.65 24.65
Claude-3.5-Sonnet 64.82 51.59 55.77
@@
Qwen-2.5-72B 80.22 69.52 95.02 40.61 8.78 20.12
Llama-3.2-3B 79.28 62.39 92.14 40.35 8.96 20.52
DeepSeek-V3 80.96 78.55 96.84 41.13 8.98 20.33
GPT-4o 80.96 77.91 96.50 40.70 8.56 19.86
Claude-3.5-Sonnet 81.17 78.90 96.77 41.13 8.99 20.00
@@
References Literature Review
[…]
Qwen-2.5-72B 40.02 28.91 33.64 38.82 29.95 9.01 15.14
Llama-3.2-3B 21.78 4.86 7.28 29.07 28.07 7.77 15.46
DeepSeek-V3 62.29 52.81 55.38 56.02 35.65 10.40 17.46
GPT-4o 60.05 50.62 54.16 59.18 30.78 9.72 15.54
Claude-3.5-Sonnet 66.43 59.06 63.06 62.32 28.59 8.90 14.41
Table 3: The experimental results of the four LLMs in Task 3. “$S_{t}$” refers to the title retrieval rate as defined in Eq 5.“Acc.” represents the accuracy defined in Eq 5. “KPR” means the Key Point Recall rate.“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
@@
Qwen-2.5-72B 7.20 12.60 13.27 12.15 17.36 18.31
Llama-3.2-3B 5.97 8.48 8.67 8.01 8.28 8.44
DeepSeek-V3 14.28 19.72 20.50 20.44 26.79 27.30
GPT-4o 8.66 18.76 19.50 18.31 27.88 28.86
Claude-3.5-Sonnet 16.06 24.34 25.21 21.20 31.90 33.25
@@
Figure 2: Distribution of LLM-generated true references over years.
@@ 4.3 Analyze LLM-Generated References from Different Dimensions
In both Task 1 and Task 3, we ask LLMs to generate references. The overall performance was discussed in the previous section. In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2. As seen in Figure LABEL:f2a, Claude-3.5 demonstrates a clear advantage over other models across all dimensions in Task 1. Additionally, the accuracy of reference generation in Task 1 for Claude-3.5, GPT-4, and Qwen-2.5 follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension. Accuracy for journal name, page, and author is also relatively high. However, DeepSeek-V3 performs worse in the author dimension compared to the other dimensions. In contrast, Llama-3.2 demonstrates higher accuracy in the page and author dimensions than in other dimensions. However, overall, Llama-3.2 does not exhibit a competitive advantage in reference generation accuracy.
Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5, DeepSeek-V3, GPT-4o, and Qwen-2.5, with particularly obvious gains in the author dimension. The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension. Notably, the accuracy of DeepSeek-V3 and GPT-4 in certain dimensions approaches or even exceeds that of Claude-3.5. However, the performance of LLaMA-3.2 remains suboptimal.
@@ 4.4 Cross-Disciplinary Analysis
We then present bar charts in Figures  LABEL:f4, which illustrate the performance of different models across various tasks and disciplines.
First, we observe that in Task 1, as shown in Figure LABEL:f4a, almost all models exhibit the highest accuracy in the Mathematics discipline and lowest accuracy in the Chemistry discipline. Additionally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences. The results are as follows: Claude-3.5 (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5 (p<.0001), Llama-3.2 (p=0.065). These results indicate significant differences in performance across disciplines for each model except Llama-3.2.
Secondly, as shown in Figure LABEL:f4b, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2. We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology. Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines. The results are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5 (p<.001), Llama-3.2 (p<.0001). These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
Additionally, we examine the accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c. It is evident that the accuracy of Claude-3.5, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5 and LLaMA-3.2 across all disciplines. Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology. The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5(p<.0001), Llama-3.2(p<.001), indicating significant differences in each models’ performance across the various disciplines.
Finally, we observe the KPR metric across different disciplines in Task 3, as shown in Figure LABEL:f4d. The results from the figure indicate that the differences between models—Claude-3.5, DeepSeek-V3, GPT-4, and Qwen-2.5—are not significant across various disciplines, a finding that is also supported by statistical tests. Specifically, the p-values are as follows: Claude-3.5 (p=0.46), DeepSeek-V3 (p=0.23), GPT-4 (p=0.18), Qwen-2.5 (p=0.10), and Llama-3.2 (p<0.001).
@@ 4.5 Comparison of LLM-Cited and Human-cited References
In this section, we also compare the LLM-generated references with those from the human-written original articles. Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article. As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5 are also cited in the human-written article, indicating a relatively high overlap rate. Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3. Notably, GPT-4o models show a 10% increase in the overlap rate. Next, we provide a more detailed comparison of the references across various dimensions. As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower. For Task 3, Claude-3.5 and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1. This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
@@ 5 Conclusion
We select four LLMs for task evaluation and find that Claude-3.5-Sonnet outperforms DeepSeek-V3, GPT-4o, Qwen-2.5-72B, and Llama-3.2-3B across all three tasks, particularly excelling in the task of generating accurate references. This advantage is likely influenced by the training data of each model. Additionally, we observed that each model has different strengths across disciplines. Overall, for the reference generation task, nearly all models perform better in Mathematics, while their performance is weaker in Chemistry and Technology. However, when writing abstracts, all models exhibit the lowest factual consistency in Social Science, as indicated by the entailment scores, compared to human-written texts.